OpenStudy (anonymous):

I asked this question an hour ago, no one helped me, I asked again here, no one is helping. Can someone just please HELP?!?!?!

OpenStudy (anonymous):

Marco and his two younger sisters would like to purchase a silver charm bracelet for their mother’s birthday. They went to the mall and found what they were looking for at Store A. In Store A, the bracelet without charms costs $85 and each charm costs $15. Use function notation that models the total price of the bracelet and how that price is based on the number of charms. Explain the reasoning behind your equation.

OpenStudy (anonymous):

This is my answer : f(x) = 85 + 15(x)

OpenStudy (anonymous):

i can try

OpenStudy (anonymous):

Just wanted to know if I'm correct

OpenStudy (anonymous):

i need to tell you one thing that im not that good in word problems but i think your right i would get other peoples opionins because im horrible at word problems

OpenStudy (anonymous):

lol thats okay :)

OpenStudy (anonymous):

im sorry but i do think you might be right im doing rates and im trying to figure those out

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

this question has different parts so.. I need help with multiple things

OpenStudy (anonymous):

What would be a reasonable domain for this function based on this scenario? Explain why this is a proper domain.

OpenStudy (anteater):

Yes, you are correct :)

OpenStudy (anonymous):

okay, but what is the domain?

OpenStudy (anteater):

Since a bracelet with no charms would be $85 and each charm adds an additional $15.

OpenStudy (anteater):

The domain would be the set of all x values you can use in your function. Since x stands for the number of charms, could you have zero charms?

OpenStudy (anonymous):

Well yeah if they choose not to have any charms.

OpenStudy (anteater):

Could you have a negative number of charms?

OpenStudy (anonymous):

No

OpenStudy (anteater):

Right, so you know that x has to be greater than or equal to 0. Is there any upper limit on the number of charms you can buy?

OpenStudy (anonymous):

So we can say that x > 0 right? would that be the domain?

OpenStudy (anteater):

x greater than or equal to 0. :) Yes, that would be the domain.

OpenStudy (anonymous):

ok thanks! but there is more..

OpenStudy (anonymous):

If Marco and his sisters have saved $250, make a graph to show all of the charms they could purchase with a bracelet if they had saved between $85 and $250.

OpenStudy (anteater):

Ok :) If you draw a graph I will give feedback on it. :)

OpenStudy (anonymous):

should it look something like this?

OpenStudy (anteater):

Yes, although you may want to change the scale of your y axis so that you can show cost up to $250

OpenStudy (anonymous):

I can't change the graph, the program wont let me but they said I can write it in words so this is how I wrote it Origin from y-axis = 85 Origin from x-axis = 0 y-axis intervals = 5 x-axis intervals = 1

OpenStudy (anonymous):

Marco found five charms that he likes. Using your function, determine the cost of the bracelet he would make. f(x) = 85 + 15(5) f(x) = 85 + 75 f(x) = 160 He would spend 160 dollars on the bracelet. Is this correct?

OpenStudy (anteater):

That labeling of your graph should work :)

OpenStudy (anonymous):

okay thanks, but there is a part 2 of this question :(

OpenStudy (anteater):

Yes, I am bouncing back and forth ... sorry! Looking at part 2 now

OpenStudy (anteater):

And yes, that is correct. Good :)

OpenStudy (anteater):

Now for the graph, what do you have?

OpenStudy (anonymous):

i couldnt make the graph cuz the program is complicated but this is the description of it.. Origin from y-axis = 85 Origin from x-axis = 0 y-axis intervals = 5 x-axis intervals = 1 gonna send you the part 2 now.. one minute

OpenStudy (anteater):

Your description of the labeling of the graph sounds good. Just a moment, please, I am also going to include a drawing.

OpenStudy (anonymous):

Marco is wondering from which store he should buy the items so that he spends the least amount of money. Which store should Marco use if he wanted the five charms and silver bracelet all together? What advice can you tell Marco to help him shop wisely depending on the number of charms he wants to buy? Justify how you reasoned your answers.

OpenStudy (anteater):

Ok. :) What do you think for this part?

OpenStudy (anonymous):

Store A f(x) = 85 + 15(5) f(x) = 85 + 75 f(x) = 160 Store B f(x) = 100 + 12(5) f(x) = 100 + 60 f(x) = 160

OpenStudy (anteater):